# Cryptanalysis of PLWE based on zero-trace quadratic roots

### v1

@@ 1. Introduction
Lattice-based cryptography has achieved an extraordinary success in the last five years, especially since the National Institute for Standards and Technology launched a public contest in 2017 to standardise different sets of quantum-resistant primitives. With four candidates proposed for standardisation, three of them belong to the lattice-based category (Crystals-Kyber for Encryption and Key Encapsulation, Crystals-Dilithium and Falcon for Digital Signatures). These suites are based on several learning problems over certain large dimensional vector spaces over finite fields, some of them with an extra ring structure.
@@
The Polynomial Learning With Errors problem (PLWE from now on) was introduced in [72]. To briefly describe it, let us denote by $R$ the quotient ring $R:=\mathbb{Z}[x]/(f(x))$ with $f(x)\in\mathbb{Z}[x]$ of degree $N$, monic and irreducible over $\mathbb{Z}[x]$, and let $q\geq 2$ be a prime. Set $R_{q}:=\mathbb{F}_{q}[x]/(f(x))$ and consider an $R_{q}$-valued uniform random distribution $U$ and an $R_{q}$-valued discrete Gaussian distribution $\chi_{R}$. The PLWE problem in decision form, the version we investigate in this work, consists in guessing, with non-negligible advantage, if a set of pairs $\{(a_{i}(x),b_{i}(x))\}_{i\geq 1}\subseteq R_{q}\times R_{q}$ of arbitrary size has been sampled either from $U\times U$ or from the PLWE distribution (see Definition 2.6 for details).
@@
In nearly all practical applications, either in post-quantum cryptography or in homomorphic encryption, the polynomial $f(x)$ defining $R_{q}$ is considered to be cyclotomic, due to the distinctive arithmetic properties enjoyed by these fields, which are used in the security reduction proofs as well as in practical implementations. However, some other works like [14], [58] and [67] deal with more general families of non-cyclotomic number fields and polynomials. Moreover, the problem is shown to enjoy a reduction from a supposedly hard problem for the family of ideal lattices whenever the prime $q$ is of the form $q=\mathcal{O}(N^{k})$ with $k\in\mathbb{N}$ independent of $N$ (for instance, see Lemma 4.1 in [72] for the power of two cyclotomic case). As a matter of fact, in most applications $k$ is considered to be $2$ and this is what we will also assume here.
@@
The aforementioned ideal-lattice based problems are supposed to be intractable in the worst case although some recent investigation shows that the Shortest Vector Problem might be more tractable for ideal lattices than for general ones (for an up-to-date account on the ideal-lattice Shortest Vector Problem in the cyclotomic case see [31], while for more general families of ideal lattices the reader is referred to the most recent [6]).
@@
Beyond worst case reductions, some particular instantiations of PLWE have been encountered to be insecure in the last few years. For instance, in [36] and [37], the authors provide an attack against PLWE whenever a simple root $\alpha\in\mathbb{F}_{q}$ of the polynomial $f(x)$ exists such that either
@@
    $\alpha$ has small residue modulo $q$.
If $f(x)=\Phi_{n}(x)$, the $n$-th cyclotomic polynomial, it is easy to check that $\alpha=\pm 1$ is never a root of $\Phi_{n}(x)$ modulo $q$ (unless $q\mid n$) and all the roots of $\Phi_{n}(x)$ have maximal order. For instance, if $q\equiv 1\pmod{n}$ so that $q$ is totally split over the ring of integers of the cyclotomic field, this order is precisely $n$. Hence cyclotomic rings are protected against conditions a) and b) above. It is however possible to attack in polynomial time the PLWE problem restricted to a large enough subring whenever a prime power degree cyclotomic polynomial splits in nonlinear factors of suitable small degree (see [10]). Further results on weak instances for the PLWE problem can be found in [57], exploiting algebraic and arithmetic vulnerabilities, and in [23],[71], exploiting a more statistical approach (the so-called Chi-square attack) which generalises the attack for case c) above.
The main contribution of the present work is to offer an extension of the attack developed by [36] to the case in which $f(x)$ has a zero-trace root $\alpha\in\mathbb{F}_{q^{k}}\setminus\mathbb{F}_{q}$ for small $k$ such that the norm $N_{\mathbb{F}_{q^{k}}|\mathbb{F}_{q}}(\alpha)$ has a suitable multiplicative order in relation with the rest of parameters in such a way that a certain look-up region is small enough. This condition is far less restricting than requiring the order to be very small.
@@
The general strategy of attacking decisional PLWE via reduction and/or trace maps to smaller spaces is already well-known. For instance, in [57, §3.2.1] it is discussed how the dual trace pairing can be used to detect whether the reduction of the error distribution modulo a prime ideal of $R$ over $q$ is detectably non-uniform; but our approach is totally different since there, the attack is based on the distribution of the values $\cos(2\pi\langle\mathbf{x},\mathbf{w}\rangle)$ with $\mathbf{x}$ belonging to certain lattice and $\mathbf{w}$ running over certain domain of the dual of that lattice.
We have structured our paper in four sections: the second recalls several definitions and notations, as well as some facts on random variables over finite fields, which we will need later on. The third is divided into two subsections: the first one recalls in detail the attack developed by [36]; motivated by the latter, the second one introduces our attack in two stages: the first stage is an attack against samples whose uniform first component belongs to a certain distinguished subring $R_{q,0}\subseteq R_{q}$ and the second stage is a general attack which reduces to the previous one in probabilistic polynomial time on the size of the set of samples and the parameters of the quotient ring. Finally, Section 4 is for simulations of our attack using Maple. The Maple code and for our examples has been located in a GitHub domain referenced therein.
@@ 2. Preliminary facts and notations
Before recalling the definition of the PLWE problem, we find convenient to explicitly mention and prove a few facts about discrete uniform and discrete Gaussian random variables over finite fields, as they are at the very background of the learning problems under study and, moreover, some of our arguments in Section 3 will involve certain manipulations of uniform distributions supported over finite dimensional vector spaces over $\mathbb{F}_{q}$.
@@
If $X_{1},X_{2},\dotsc,X_{N}$ are independent uniform distributions over $\mathbb{F}_{q}$ then, for each $\lambda_{1},\lambda_{2},\dotsc,\lambda_{N}\in\mathbb{F}_{q}$, not all of them zero, the variable $\sum_{i=1}^{N}\lambda_{i}X_{i}$ is uniform.
Proof.
First, it is obvious that given a uniform distribution $X$ over $\mathbb{F}_{q}$ and given $\lambda\in\mathbb{F}_{q}^{*}$, the variable $\lambda X$ is uniform. So, it is enough to prove that if $X_{1}$ and $X_{2}$ are uniform, then the variable $X_{1}+X_{2}$ is uniform. But for $i\in\mathbb{F}_{q}$, using the Total Probability Theorem we have
$P[X_{1}+X_{2}=i]=\sum_{j\in\mathbb{F}_{q}}P[X_{1}+X_{2}=i|X_{2}=j]P[X_{2}=j]\\ =q^{-1}\sum_{j\in\mathbb{F}_{q}}P[X_{1}=i-j]=\frac{1}{q}.$
@@
Let $X$ be a uniform distribution on the vector space $\mathbb{F}_{q}^{N}$ and let $A:\mathbb{F}_{q}^{N}\to\mathbb{F}_{q}^{M}$ be a linear map. Then, the distribution $AX$ obtained by a) sampling an element $\mathbf{x}\in\mathbb{F}_{q}^{N}$ uniformly from $X$ and b) returning $A\mathbf{x}$ is also uniform over $\mathrm{Im}(A)$. In particular, the $i$-th component is uniformly sampled from $\mathbb{F}_{q}$.
@@
The second statistical component of our learning problem is the notion of a discrete Gaussian random variable. Here we follow closely [36, Section 2] and [47, Section 2].
As it is well known, for $r>0$, the Gaussian function $\rho_{r}:\mathbb{R}^{N}\to\mathbb{R}_{>0}$ defined by $\rho_{r}(\mathbf{x})=\exp(-\pi||\mathbf{x}||^{2}/r^{2})$ yields, after normalising, the density function of a continuous Gaussian random variable of zero mean and variance-covariance matrix $r^{2}I_{N}$, where $I_{N}$ is the $N$-dimensional identity matrix. Furthermore, fixing a basis $\{\mathbf{h}_{i}\}_{i=1}^{N}$ of $\mathbb{R}^{N}$ and given a vector $\mathbf{r}=(r_{1},\dotsc,r_{N})\in\mathbb{R}_{+}^{N}$, if $D_{r_{i}}$ is a $1$-dimensional continuous Gaussian random variable of zero mean and variance $r_{i}^{2}$ and if these $N$ random variables are independent, then the variable $D_{\mathbf{r}}=\sum_{i=1}^{N}D_{r_{i}}\mathbf{h}_{i}$ is a continuous elliptic $N$-dimensional Gaussian variable of zero mean and whose variance-covariance matrix has $\mathbf{r}$ as main diagonal and 00 elsewhere. Denote by $\rho_{\mathbf{r}}(\mathbf{x})$ the density function of $D_{\mathbf{r}}$.
@@
$P[X=\mathbf{x}]=\frac{\rho_{\mathbf{r}}(\mathbf{x})}{\rho_{\mathbf{r}}(\mathcal{L})}\quad\text{for any }\mathbf{x}\in\mathcal{L}.$
@@ 2.1. The PLWE problem
We can naturally endow the ring $R$ with a lattice structure over $\mathbb{R}^{N}$ via the coefficient embedding:
@@
Now, for a prime $q\geq 2$, if $\chi$ is a discrete Gaussian random variable with values in $R$ (or, rather, in its image $\sigma_{coef}(R)\subseteq\mathbb{R}^{N}$ by the coefficient embedding) we can also reduce its outputs modulo $q$. This is hence a random variable with finite support and we refer to it as a discrete Gaussian variable modulo $q$.
Let $q\in\mathbb{Z}$ be a fixed prime and $\chi_{R}$ be the reduction modulo $q$ of a discrete $R$-valued Gaussian random variable of 00 mean and variance-covariance matrix $\Sigma_{R}$, hence with values in $R_{q}$.
@@
For a fixed element $s\in R_{q}$ and an error distribution $\chi_{R}$ over $R_{q}$, the PLWE distribution attached to the triple $(R_{q},s,\chi_{R})$ is the probabilistic algorithm $\mathcal{A}_{s,\chi_{R}}$ that works as follows:
@@
    Samples an element $e\in R_{q}$ from $\chi_{R}$.
@@
Denoting by $R^{2}_{q}$ the Cartesian product $R_{q}\times R_{q}$, the PLWE problem for the triple $(R,q,\chi_{R})$ is defined as follows. We give the decision version, since that is all we need for our purpose.
Definition 2.7 (PLWE decision problem).
Let $\chi_{R}$ be defined as before. The PLWE problem consists in deciding with non-negligible advantage, for a set of samples of arbitrary size $(a_{i},b_{i})\in R_{q}^{2}$, whether they are sampled from the PLWE distribution or from the uniform distribution.
Notice that the ring $R$ is embedded into $\mathbb{R}^{N}$ by the $\sigma_{coef}$. In particular, our discrete Gaussian distribution $\chi_{R}$ takes values over the quotient ring $R_{q}$, as in [36] and [37].
@@ 3. An attack based on traces over finite extensions of $\mathbb{F}_{q}$
The first part of this Section will present in detail the attack developed in [36], since our proposal is an extension of such attack to the case in which $f(x)$ has a zero-trace root $\alpha\in\mathbb{F}_{q^{k}}\setminus\mathbb{F}_{q}$. Our attack will be then described in the second part of the section.
@@ 3.1. Attack developed by [36] and [37]
    $\alpha$ has small residue modulo $q$.
The expressions small order and small residue are not made fully precise herein, but the examples provided by the authors show that by such a term we must understand orders of up to $5$ while by small residue the authors understand $\alpha=2,3$.
@@
Assume that $\alpha=1$ is a root of $f(x)$. Given a PLWE sample $(a(x),b(x)=a(x)s(x)+e(x))$, the error term, $e(x)=\sum_{i=0}^{N-1}e_{i}x^{i}$, has its coefficients $e_{i}\in\mathbb{F}_{q}$ sampled from a discrete Gaussian variable of small enough standard deviation $\sigma$ (the authors consider ranges for $\sigma$ in the interval $[0.3,8]$, as suggested by the applications they analyse). For an element $s\in\mathbb{F}_{q}$, writing $s=s(1)$ and applying the evaluation map, we have
$b(1)-a(1)s=e(1)=\sum_{i=0}^{N-1}e_{i},$
and the sum $\sum_{i=0}^{N-1}e_{i}$ is hence sampled from a discrete Gaussian variable of standard deviation $\sqrt{N}\sigma$, which according to practical specifications, is of order $O(q^{1/4})$.
For a right guess $s=s(1)$, the value $b(1)-a(1)s$ will belong to the set of integers $[-\sqrt{N}\sigma,\sqrt{N}\sigma]\cap\mathbb{Z}$ (which can be easily enumerated) with probability at least $0.65$. Hence, we will refer to the set $[-\sqrt{N}\sigma,\sqrt{N}\sigma]\cap\mathbb{Z}$ as the smallness region for this attack.
In the same paper, the authors develop a more subtle attack, based on the multiplicative order of the roots of $f(x)$. Suppose that $r$ is the order of a root $\alpha\in\mathbb{F}_{q},\alpha\neq 1$. Again given a PLWE sample $(a(x),b(x)=a(x)s(x)+e(x))$, for $s=s(\alpha)\in\mathbb{F}_{q}$, we have
$b(\alpha)-a(\alpha)s=e(\alpha)=\sum_{j=0}^{r-1}\sum_{i=0}^{\left\lfloor\frac{N}{r}\right\rfloor-1}e_{ir+j}\alpha^{j}.$
Now the terms $e_{j}=\sum_{i=0}^{\left\lfloor\frac{N}{r}\right\rfloor-1}e_{ir+j}$ come from a discrete Gaussian distribution of 00 mean and standard deviation less than or equal $\sqrt{N/r}\sigma$, hence they will belong to the set of integers $[-\sqrt{N/r}\sigma,\sqrt{N/r}\sigma]\cap\mathbb{Z}$ with probability at least $0.65$. This leads us to consider in this case the smallness region as the set $\Sigma$ of all possible values for $e(\alpha)$, which can be precomputed and stored in a look-up table. Observe that
$|\Sigma|\leq\left(2\sqrt{N/r}\sigma+1\right)^{r}.$

### v3

@@ Abstract.
Key words and phrases:
Lattice-based cryptography, algebraic cryptanalysis, orders in finite fields, traces over finite fields
2010 Mathematics Subject Classification:
11T71, 94A60
@@ 1. Introduction
Lattice-based cryptography has achieved an extraordinary success in the last years, especially since the National Institute for Standards and Technology launched a public contest in 2017 to standardise different sets of quantum-resistant primitives. In 2022, four candidates were proposed for standardisation, three of them belonging to the lattice-based category (Crystals-Kyber for Encryption and Key Encapsulation, Crystals-Dilithium and Falcon for Digital Signatures). These suites are based on several learning problems over certain large-dimensional vector spaces over finite fields, some of them with an extra ring structure.
@@
In the present work, we are concerned in particular with the Polynomial Learning With Errors problem (PLWE from now on), which was introduced in [23]. To briefly describe it, let us denote by $R$ the quotient ring $R:=\mathbb{Z}[x]/(f(x))$ with $f(x)\in\mathbb{Z}[x]$ of degree $N$, monic and irreducible over $\mathbb{Z}[x]$, and let $q\geq 2$ be a prime. Set $R_{q}:=\mathbb{F}_{q}[x]/(f(x))$ and consider an $R_{q}$-valued uniform random distribution $U$ and an $R_{q}$-valued discretized Gaussian distribution $\chi_{R_{q}}$. The PLWE problem in decision form, the version we investigate in this work, consists in guessing, with non-negligible advantage, if a set of pairs $\{(a_{i}(x),b_{i}(x))\}_{i\geq 1}\subseteq R_{q}\times R_{q}$ of arbitrary size has been sampled either from $U\times U$ or from the PLWE distribution (see Definition 2.6 for details).
In nearly all practical instances of (P)LWE-based cryptographic schemes, either in post-quantum cryptography or in homomorphic encryption, the polynomial $f(x)$ defining $R_{q}$ is considered to be cyclotomic, due to the distinctive arithmetic properties enjoyed by these fields, which are used in the security reduction proofs as well as in practical implementations. For instance, the NIST standard Crystals-Kyber ([17]) is based on the Module Learning With Errors problem, a variant of PLWE. In particular, its modulus polynomial is the $2^{8}$ cyclotomic polynomial. The third-round NIST candidate SABER ([4]) also uses a variant of PLWE backed on the $2^{8}$-th cyclotomic polynomial. As for homomorphic encryption standards, nearly all the schemes use PLWE moduli polynomials that are either cyclotomic (mainly HELib [1], LATTIGO [2], or OpenFHE [3]) or minimal polynomials of the maximal totally real extension (the current implementation of LATTIGO allows to choose this subfield). For a complete analysis of homomorphic encryption schemes based on non-cyclotomic number fields and their practical performance we refer the reader to [18].
@@
However, some other works like [20] and [21] deal with more general families of non-cyclotomic number fields and polynomials. Moreover, the problem is shown to enjoy a reduction from a supposedly hard problem for the family of ideal lattices whenever the prime $q$ is of the form $q=\mathcal{O}(N^{k})$ with $k\in\mathbb{N}$ independent of $N$ (for instance, see Lemma 4.1 in [23] for the power-of-two cyclotomic case). As a matter of fact, in most applications $k$ is considered to be $2$ and this is what we will also assume here.
@@
The aforementioned ideal-lattice based problems are supposed to be intractable in the worst case although some further investigation shows that the Shortest Vector Problem might be more tractable for ideal lattices than for general ones (for an up-to-date account on the ideal-lattice Shortest Vector Problem in the cyclotomic case see [11], while for more general families of ideal lattices the reader is referred to [5]).
@@
Beyond worst-case reductions, some particular instantiations of PLWE have been encountered to be insecure. For instance, in [13] and [14], the authors provide an attack against PLWE whenever a simple root $\alpha\in\mathbb{F}_{q}$ of the polynomial $f(x)$ exists such that either
@@
    $\alpha$ has small residue modulo $q$, or
  4. d)
    none of the above situations holds but $\alpha$ and the variance of the error distribution are such that one can distinguish the PLWE distribution from the uniform one.
If $f(x)=\Phi_{n}(x)$, the $n$-th cyclotomic polynomial, it is easy to check that $\alpha=\pm 1$ is never a root of $\Phi_{n}(x)$ modulo $q$ (unless $q\mid n$) and all the roots of $\Phi_{n}(x)$ have maximal order. For instance, if $q\equiv 1\pmod{n}$ so that $q$ is totally split over the ring of integers of the cyclotomic field, this order is precisely $n$. Hence any attack based on conditions a) or b) does not apply to cyclotomic rings. It is however possible to attack in polynomial time the PLWE problem restricted to a large enough subring whenever a prime-power-degree cyclotomic polynomial splits in nonlinear factors of suitable small degree (see [8]). This earlier attack addresses a situation not dealt with in [13], which assumes that the modulus polynomial is totally split, but the assumption that the samples must belong to a smaller subring imposes a strong limitation on the attack to work.
The family of attacks based on case d) has been extended in [6], [7] by the second and third author, jointly with R. Martín Sánchez-Ledesma. In that recent work, new sets of dangerous instances have been identified. In addition, [6] elaborates on the results of the present work and provides new attacks when the modulus polynomial admits a cubic or quartic irreducible factor, whose roots have zero trace and a norm with small order. However, even if [6] has been published first, the present work provides foundational background which is omitted and taken for granted in [6]. In particular, in the present work we prove Proposition 3.5, which provide the complexity computation of our main algorithm. Likewise, we prove Lemma 3.6 and Lemma 3.7, which reduce the analysis of traces of errors evaluated on roots in $\mathbb{F}_{q^{2}}$ to the identification of a smallness region in $\mathbb{F}_{q}$.
Further results on weak instances for the PLWE problem can be found in [19], exploiting algebraic and arithmetic vulnerabilities, and in [10],[22], exploiting a more statistical approach (the so-called Chi-square attack) which generalises the attack for the case c) above.
The main contribution of the present work is to offer an extension of the attack developed by [13] to the case in which $f(x)$ has a zero-trace root $\alpha\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ such that the norm $N_{\mathbb{F}_{q^{2}}\mid\mathbb{F}_{q}}(\alpha)$ has a suitable multiplicative order in relation with the rest of parameters in such a way that the trace of the error polynomials evaluated at $\alpha$ belong to a region whose cardinality is small enough in relation with $q$. This condition is far less restricting than requiring the order of the norm to be very small.
@@
The general strategy of attacking the decision PLWE problem via reduction and/or trace maps to smaller spaces is already well known. For instance, in [19, §3.2.1] it is discussed how the dual trace pairing can be used to detect whether the reduction of the error distribution modulo a prime ideal of $R$ over $q$ is uniform or not; but our approach is totally different, since we do not resort to dual lattices at all.
We have structured our paper in four sections: Section 2 recalls several definitions and notations, as well as some facts on random variables over finite fields, which we will need later on. Section 3 consists of two subsections: the first one recalls in detail the attack developed by [13]; motivated by the latter, the second one introduces our attack in two stages: the first stage is an attack against samples whose uniform first component belongs to a certain distinguished subring $R_{q,0}\subseteq R_{q}$ and the second stage is a general attack which reduces to the previous one in probabilistic polynomial time on the size of the set of samples and the parameters of the quotient ring. Finally, Section 4 describes simulations of our attack using Maple. The Maple code for our examples has been uploaded to a GitHub domain referenced therein.
Finally, we must stress out that neither this work nor [13], [12], [19], [21] does not attempt to recommend or rule out any precise sets of parameters for PLWE-based cryptosystems. Our aim is just to provide a further study of algebraic cryptanalysis towards these schemes, which is to be regarded as a natural continuation of the aforementioned works and shows how the trade-off between the order of the traces modulo $q$, the variance and the dimension impact on the security of the cryptosystem.
@@ 2. Preliminary facts and notations
Before recalling the definition of the PLWE problem, we find it convenient to explicitly mention and prove a few facts about discrete uniform and discrete Gaussian random variables over finite fields, as they are at the very background of the learning problems under study and, moreover, some of our arguments in Section 3 will involve certain manipulations of uniform distributions supported over finite dimensional vector spaces over $\mathbb{F}_{q}$.
@@
([8, Lemma 2.2]) If $X_{1},X_{2},\dotsc,X_{N}$ are independent uniform distributions over $\mathbb{F}_{q}$ then, for each $\lambda_{1},\lambda_{2},\dotsc,\lambda_{N}\in\mathbb{F}_{q}$, not all of them zero, the variable $\sum_{i=1}^{N}\lambda_{i}X_{i}$ is uniform.
@@
Let $X$ be a uniform distribution on the vector space $\mathbb{F}_{q}^{N}$ and let $A:\mathbb{F}_{q}^{N}\to\mathbb{F}_{q}^{M}$ be a linear map. Then, the distribution $AX$ obtained by a) sampling an element $\mathbf{x}\in\mathbb{F}_{q}^{N}$ uniformly from $X$ and b) returning $A\mathbf{x}$, is also uniform over $\mathrm{Im}(A)$. In particular, the $i$-th component is uniformly sampled from $\mathbb{F}_{q}$.
@@
The second statistical component of our learning problem is the notion of a discrete Gaussian random variable. Here we follow closely [13, Section 2] and [16, Section 2].
[…]
@@
$P[X=\mathbf{x}]=\frac{\rho_{\mathbf{r}}(\mathbf{x})}{\rho_{\mathbf{r}}(\mathcal{L})}\quad\text{for any }\mathbf{x}\in\mathcal{L}\text{ where }\rho_{\mathbf{r}}(\mathcal{L})=\sum_{\mathbf{x}\in\mathcal{L}}\rho_{\mathbf{r}}(\mathbf{x}).$
@@ 2.1. The PLWE problem
We can naturally endow the ring $R$ with a lattice structure over $\mathbb{R}^{N}$ via the following group monomorphism known in the literature as the coefficient embedding:
@@
For a prime $q\geq 2$, if $\chi_{R}$ is a discrete Gaussian random variable with values in $R$ (or, rather, in its image $\sigma_{coef}(R)\subseteq\mathbb{R}^{N}$ by the coefficient embedding) we can also reduce its outputs modulo $q$. This is hence a random variable with finite support and we refer to it as a discrete Gaussian variable modulo $q$ and denote it by $\chi_{R_{q}}$.
Now, let $q\in\mathbb{Z}$ be a fixed prime and let $\chi_{R_{q}}$ be such a $q$-reduced discrete Gaussian random variable of 0 mean, hence with values in $R_{q}$. Then we have the following
@@
For a fixed element $s\in R_{q}$ and an error distribution $\chi_{R_{q}}$ over $R_{q}$ the PLWE distribution attached to the triple $(R_{q},s,\chi_{R_{q}})$ is the probabilistic algorithm $\mathcal{A}_{s,\chi_{R_{q}}}$ that works as follows:
@@
    Samples an element $e\in R_{q}$ from $\chi_{R_{q}}$.
@@
Denoting by $R^{2}_{q}$ the Cartesian product $R_{q}\times R_{q}$, the PLWE problem for the triple $(R,q,\chi_{R_{q}})$ is defined as follows. We give the decision version, since that is all we need for our purpose.
Definition 2.7 (PLWE decisional problem).
Let $\chi_{R_{q}}$ be defined as before. Let $\mathcal{A}$ be the PLWE distribution attached to a triple $(R_{q},s,\chi_{R_{q}})$ for some $s\in R_{q}$ and $\mathcal{B}$ the uniform distribution on $R_{q}^{2}$. The (decisional) PLWE problem consists in deciding with non-negligible advantage, for a set of samples of arbitrary size $(a_{i},b_{i})\in R_{q}^{2}$, whether they are sampled from $\mathcal{A}$ or from $\mathcal{B}$.
@@ 3. An attack based on traces over finite extensions of $\mathbb{F}_{q}$
The first part of this Section will present in detail the attack developed in [13], since our proposal is an extension of such attack to the case in which $f(x)$ has a zero-trace root $\alpha\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$. Our attack will be then described in the second part of the section.
We point out that although not explicitly stated in Definition 2.7 of the PLWE problem, it is assumed that the error distribution is a zero-mean discrete spherical $q$-reduced Gaussian, $\chi_{R_{q}}$. Remark that most security reduction proofs actually assume that $\chi_{R_{q}}$ is spherical: this assumption is legitimate as long as only a uniform upper-bound on the diagonal entries is required, as is the case in most of these security reduction proofs and also in ours.
Moreover, we will assume, as in [13], that $\chi_{R_{q}}$ is truncated at $2\sigma$, namely, that each of the independent Gaussian distributions corresponding to each coordinate is so truncated and renormalized. In particular, the support of each coordinate’s probability mass function -is contained in $2\sigma,2\sigma]\cap\mathbb{Z}$.
@@ 3.1. Attack developed by [13] and [14]
    $\alpha$ has small residue modulo $q$, or
  4. d)
    none of the above situations holds but $\alpha$ and the variance are such that one can distinguish the PLWE distribution from the uniform one.
The expressions small order and small residue are not made fully precise therein, but the examples provided by the authors show that by such a term we must understand orders of up to $5$ while by small residue the authors understand $\alpha=2,3$. We will not address cases c) and d), which are investigated in [6].
@@
Assume that $\alpha=1$ is a root of $f(x)$. Given a PLWE sample $(a(x),b(x)=a(x)s(x)+e(x))$, the error term, $e(x)=\sum_{i=0}^{N-1}e_{i}x^{i}$, has its coefficients $e_{i}\in\mathbb{F}_{q}$ sampled from $\chi_{R_{q}}$ of small enough standard deviation $\sigma$ (the authors consider ranges for $\sigma$ in the interval $[0.3,8]$, as suggested by the applications they analyse). Applying the evaluation map, we have
$b(1)-a(1)s(1)=e(1)=\sum_{i=0}^{N-1}e_{i},$
and the sum $\sum_{i=0}^{N-1}e_{i}$ hence takes values on the set $2N\sigma,2N\sigma]\cap\mathbb{Z}$.
If one loops over $\mathbb{F}_{q}$ for a guess for $s(1)$, a right guess $g$ is one such that the value b(1)a(1)gb(1)-a(1)g belongs to $\Sigma$. Since the error distribution is assumed to be a centered Gaussian truncated at $2\sigma$, we can detect which distribution do the samples come from (uniform or PLWE) with significant probability. We will refer to the set $\Sigma$ as the smallness region for this attack.
In the same paper, the authors develop a more subtle attack, based on the multiplicative order of the roots of $f(x)$. Suppose that $r$ is the order of a root $\alpha\in\mathbb{F}_{q},\alpha\neq 1$. Again given a PLWE sample $(a(x),b(x)=a(x)s(x)+e(x))$, we have
$b(\alpha)-a(\alpha)s(\alpha)=e(\alpha)=\sum_{j=0}^{r-1}\sum_{i=0}^{\left\lfloor\frac{N}{r}\right\rfloor-1}e_{ir+j}\alpha^{j}.$
Now each coefficient $\epsilon_{j}:=\sum_{i=0}^{\left\lfloor\frac{N}{r}\right\rfloor-1}e_{ir+j}$ takes values on $2\frac{N}{r}\sigma,2\frac{N}{r}\sigma]\cap\mathbb{Z}$. Hence, in this case we define the smallness region:
$\Sigma:=\left\{\sum_{j=0}^{r-1}x_{j}\alpha^{j}:\,\,x_{j}\in\left[-2\frac{N}{r}\sigma,2\frac{N}{r}\sigma\right]\cap\mathbb{Z}\right\},$
namely, the set of all possibilities for $e(\alpha)$. $\Sigma$ can be precomputed and stored as a look-up table.
Observe that
$|\Sigma|\leq\left(4\frac{N}{r}\sigma+1\right)^{r}.$
@@
  • set $G:=\emptyset$
  • for $g\in\mathbb{F}_{q}$ do
    • for $(a_{i}(x),b_{i}(x))\in C$ do
        if $b_{i}(\alpha)-a_{i}(\alpha)g\notin\Sigma$ then
          next $g$
    • set $G:=G\cup\{g\}$
  • if $G=\emptyset$ then return NOT PLWE
  • if $G=\{g\}$ then return $g$
  • if $|G|>1$ then return NOT ENOUGH SAMPLES